# Learned Image Transmission with Hierarchical Variational Autoencoder

### v2

@@ Abstract
In this paper, we introduce an innovative hierarchical joint source-channel coding (HJSCC) framework for image transmission, utilizing a hierarchical variational autoencoder (VAE). Our approach leverages a combination of bottom-up and top-down paths at the transmitter to autoregressively generate multiple hierarchical representations of the original image. These representations are then directly mapped to channel symbols for transmission by the JSCC encoder. We extend this framework to scenarios with a feedback link, modeling transmission over a noisy channel as a probabilistic sampling process and deriving a novel generative formulation for JSCC with feedback. Compared with existing approaches, our proposed HJSCC provides enhanced adaptability by dynamically adjusting transmission bandwidth, encoding these representations into varying amounts of channel symbols. Additionally, we introduce a rate attention module to guide the JSCC encoder in optimizing its encoding strategy based on prior information. Extensive experiments on images of varying resolutions demonstrate that our proposed model outperforms existing baselines in rate-distortion performance and maintains robustness against channel noise. The source code will be made available upon acceptance.
Index Terms:
Deep learning, image transmission, joint source-channel coding, semantic communication.
I Introduction
@@
To meet the transmission requirements of heavy data traffic in future sixth-generation (6G) networks, wireless edge devices need to be equipped with higher transmission efficiency. Most contemporary systems employ a two-step strategy for data transmission: first, the raw data is compressed using a source codec, such as JPEG [1] and BPG [2]. Then, the encoded bits are protected with redundancy introduced by a carefully designed channel codec, such as LDPC and Polar codes [3]. However, in many practical applications, the bit length is generally finite, making it impossible to guarantee optimality. In this context, joint source-channel coding (JSCC) has emerged as a potential solution, offering higher coding gains than the traditional separation-based coding paradigm.
@@
With the revolutionary progress of deep learning in various fields, such as image compression [4, 5, 6] and generative models [7, 8], a novel design paradigm for JSCC, called learned image transmission (LIT), has been conceived by formulating the communication pipeline as an end-to-end deep learning model [9, 10, 11], as known as semantic communication. Specifically, these methods leverage powerful neural networks to implement the encoding and decoding processes. In this approach, the whole system is viewed as an autoencoder (AE), which can be jointly learned in a data-driven manner. A notable method proposed by [9] employed CNNs to construct the source and channel codecs for wireless image transmission, achieving great performance by mapping the input image directly into channel symbols. Moreover, the authors of [12, 13] investigate the JSCC with feedback. In this context, the transmission of these representations is divided into multiple phases, with the transmitter receiving the channel symbol vector after each phase, which simplifies the encoding process and improves the overall performance.
@@
Beyond deterministic AEs, some studies have employed variational autoencoders (VAE) to design JSCC systems [14, 15, 16], where the channel symbols are generated through sampling. Part of these VAE-based methods show superior performance compared to deterministic AE-based methods, particularly under severe channel conditions. Though VAE-based methods have demonstrated remarkable performance, they experience significant performance degradation on high-resolution images. Furthermore, most existing methods only support fixed-rate coding, which contrasts with emerging works on transform coding-based image compression [4], where the compression rate for each image is determined by the estimated entropy of its feature representation and varies with different samples. Consequently, these methods are less flexible and adaptive, potentially leading to performance penalties.
@@
In this work, we aim to overcome the limitations of previous methods while enhancing performance. Specifically, we develop a hierarchical JSCC (HJSCC) framework based on a powerful hierarchical VAE architecture [17]. Our transmitter employs both bottom-up and top-down paths to autoregressively generate multiple hierarchical representations of the original image. These representations are then mapped to channel symbols using multiple JSCC encoder blocks. Building upon this, we further explore the application of HJSCC in a classical scenario where a feedback link exists. By modeling transmission over a noisy channel as a probabilistic sampling process, we derive a novel generative formulation for JSCC with feedback, which achieves significantly better performance than most existing advanced schemes. While there have been attempts at variable-rate transmission [18, 19] in the realm of JSCC without feedback, the problem of rate-adaptive design for JSCC with feedback remains underexplored. Unlike existing works [12, 13], we leverage the prior distribution (which characterizes the entropy information) of each representation to generate masks that control the number of symbols for each representation. This approach allows us to dynamically adjust the transmission rate. Additionally, we introduce a rate attention module to guide the JSCC encoder in adjusting the encoding strategy according to its prior information.
@@ II Related Works
VAE can be employed as deep generative models capable of generating high-dimensional data based on a low-dimensional latent space, and is a variant of autoencoder [20, 21]. By sampling from the learned latent space distribution and passing these samples through the decoder network, VAE can generate new data points that resemble the training data. However, the original VAE is known to perform worse than many other generative models, particularly when applied to high-resolution images. [22, 23, 24] addressed this by proposing a deep hierarchical VAE, where the latent variable is divided into several disjoint groups, achieving significantly better performance than standard VAEs. More recently, VAE has been applied to compression tasks [25, 26, 27, 28].
@@
Unlike the separation-based design described above, recent studies have delved into the utilization of AE and its variants, e.g., VAE to design wireless image transmission systems, resulting in a number of efficient methods [9, 29, 10, 11, 29, 30, 31]. In particular, [9, 32] and [30] conceived of using neural networks to simultaneously finish the source encoding/decoding and channel coding/decoding, with the goal of jointly optimizing the entire system to maximize PSNR. The VAE-based methods adopted probabilistic modeling, wherein the encoding process is characterized as a stochastic procedure. In these systems, channel symbols are generated by sampling from the probability distribution conditioned on the input image [15, 31]. These approaches have demonstrated superior performance, especially under severe transmission conditions.
III Proposed Methods
III-A Background
@@
III-A1 VAE and Hierarchical VAE
As stated in [21], VAE is a stochastic variational inference scheme that can be applied to various intelligent tasks, such as recognition, denoising, and generation. As shown in Fig. 1(a), to formulate a vision-related model, we typically start with the following premises. Let $\mathbf{x}$ denote an image intensity vector, which is drawn from a dataset $\mathcal{X}$ with distribution $p_{\mathbf{x}}$. Another variable is the latent variable $\mathbf{z}$, with a prior $p_{\mathbf{z}}$. The main target of a VAE is to learn a generative model (decoder) $p_{\mathbf{x}|\mathbf{z}}$ for sampling, and a posterior density model (encoder) $q_{\mathbf{z}|\mathbf{x}}$ for variational inference. The objective for learning a VAE model can be formulated as minimizing the (variational) upper bound on the marginal likelihood of a batch of data points, as given by
$\mathcal{L}\left(\bm{\theta},\bm{\phi}\right)=\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\big{[}D_{KL}\left(q_{\mathbf{z}|\mathbf{x}}\|p_{\mathbf{z}}\right)-\log p_{\mathbf{x}|\mathbf{z}}\big{]},$ (1)
where $D_{KL}\left(q_{\mathbf{z}|\mathbf{x}}\|p_{\mathbf{z}}\right)=\log\big{(}q_{\mathbf{z}|\mathbf{x}}/p_{\mathbf{z}}\big{)}$, $\bm{\theta}$ and $\bm{\phi}$ represent the parameters of the encoder $q_{\mathbf{z}|\mathbf{x}}$ and decoder $p_{\mathbf{x}|\mathbf{z}}$, respectively.
@@ III-A1 VAE and Hierarchical VAE
$\mathcal{L}\left(\bm{\theta},\bm{\phi}\right)=\quad\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\left[\sum_{l=1}^{L}D_{KL}\left(q_{\mathbf{z}_{l}|\mathbf{z}_{<l},\mathbf{x}}\|p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}\right)\!-\!\log p_{\mathbf{x}|\mathbf{z}}\right],$ (2)
@@
III-B Proposed HJSCC
III-B1 System Overview
@@
Here, we aim to give a brief overview of deep learning-based JSCC. In particular, the model of a typical JSCC is shown in Fig. 2. The transmitter employs a JSCC encoder to map the input image $\mathbf{x}\in\mathbb{R}^{N}$ directly into channel symbol vector $\mathbf{s}\in\mathbb{C}^{K}$ for transmission, where $N$ denotes the number of pixels and $K$ represents the number of channel symbols. This process can be expressed as $\mathbf{s}=g_{e}(\mathbf{x};\bm{\theta}_{e})$, where $g_{e}$ signifies the encoding function and $\bm{\theta}_{e}$ represents its trainable parameters. Accounting for the limited transmission power, the transmitted signal should satisfy the power constraint $P$, implying that $\|\mathbf{s}\|^{2}_{2}/K\leq P$. Subsequently, the channel symbol vector is transmitted through the wireless channel, as given by $\mathbf{\tilde{s}}=\mathbf{s}+\mathbf{n}$, where $\mathbf{n}\sim\mathcal{N}(0,\sigma^{2}_{n}\mathbf{I})$ is additive white Gaussian noise (AWGN). At the receiver, the received noisy signal $\mathbf{\tilde{s}}$ is processed by a decoder $g_{d}$ to obtain the reconstructed image $\mathbf{\bar{x}}=g_{d}(\mathbf{\tilde{s}};\bm{\phi}_{d})$, with $\bm{\phi}_{d}$ denoting the trainable parameters of the JSCC decoder. The optimization objective of a deep learning-based JSCC system is to minimize the difference between $\mathbf{x}$ and $\mathbf{\bar{x}}$, and thus mean-square error (MSE) can be employed as the loss function. Furthermore, to evaluate the performance of a JSCC system and ensure fairness, we define the signal-to-noise ratio (SNR) as SNR =$\text{SNR }=10\log\frac{P}{\sigma^{2}}(\mathrm{dB})$, which characterizes the channel quality of the system. Then, we introduce the channel bandwidth ratio (CBR) to describe the transmission rate (overhead), which is expressed as $\text{CBR}=K/N$. Intuitively, CBR actually signifies the number of symbols for transmitting one pixel, and a higher CBR brings a higher overhead, while usually resulting in a better system performance.
@@
While previous studies have achieved excellent rate-distortion performance based on the framework depicted in Fig. 2, it is evident that the transmission rate is solely determined by the image resolution. This limitation can result in performance degradation in overall rate-distortion due to the inability to adaptively adjust the rate for each image [33]. To address this issue, we propose our HJSCC framework, as illustrated in Fig. 3. Given an image $\mathbf{x}$, the bottom-up path generates a set of latent features, which are subsequently passed to the top-down path to autoregressively generate the latent representations $\bm{\mu}\triangleq\left\{\bm{\mu}_{1},\bm{\mu}_{2},\ldots,\bm{\mu}_{L}\right\}$. These representations are then fed to a set of JSCC encoders $g_{e}=\left\{g^{1}_{e},g^{2}_{e},\ldots,g_{e}^{L}\right\}$, respectively. In this way, we are able to obtain a set of channel symbol vectors $\mathbf{s}\triangleq\left\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L}\right\}$, where $\mathbf{s}_{l}=g_{e}^{l}(\mathbf{z}_{l})$. Then, these channel symbol vectors are transmitted to the receiver through the wireless link, and the received symbol vectors are represented by $\mathbf{\tilde{s}}\triangleq\left\{\mathbf{\tilde{s}}_{1},\mathbf{\tilde{s}}_{2},\ldots,\mathbf{\tilde{s}}_{L}\right\}$. At the receiver, the noisy $\mathbf{\tilde{s}}$ undergoes processing by the JSCC decoder $g_{d}^{l}$, and we obtain $\bm{\tilde{\mu}}\triangleq\left\{\bm{\tilde{\mu}}_{1},\bm{\tilde{\mu}}_{2},\ldots,\bm{\tilde{\mu}}_{L}\right\}$. With $\bm{\tilde{\mu}}$ at hand, the receiver can reconstruct the image using the top-down path (decoder), and the reconstructed image is denoted by $\mathbf{\hat{x}}_{\text{H}}$. Our objective is to minimize the distortion between the transmitted image $\mathbf{x}$ and $\mathbf{\hat{x}}_{\text{H}}$.
@@
III-C Training Objective Formulation
We aim to develop a rate-adaptive JSCC model capable of adjusting the transmission rate, CBR, based on the source content. To this end, we propose an inherited training strategy motivated by [18], where the objective of minimizing $d(\mathbf{x},\mathbf{\hat{x}}_{\text{H}})$ is guided by a learned image coder. Specifically, we have designs for the posteriors, priors, and training objectives as follows [25].
@@ III-C Training Objective Formulation
where $\mathcal{U}$ denotes a uniform distribution centered on $\mu_{l}^{(i)}$, and $\mu_{l}^{(i)}$ is the $i$-th element of parameter $\bm{\mu}_{l}$, which is obtained from the $l$-th posterior branch.
Priors: The conditional prior distribution $p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$ is defined to be a Gaussian distribution convolved with a uniform distribution [4]:
@@
Training Loss: With the posteriors and priors defined above, the loss function (2) for an image compression model [25] can be expressed as
@@
$\mathcal{L}=\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\big{[}\underbrace{\sum_{l=1}^{L}-\alpha\log p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}(\mathbf{z}_{l}|\mathbf{z}_{<l})}_{\text{transmission rate}}+\lambda\cdot(\underbrace{d(\mathbf{x},\mathbf{\hat{x}})+d(\mathbf{x},\mathbf{\hat{x}}_{\text{H}}))}_{\text{weighted distortion}}\big{]},$ (6)
where the calculation ways of each variable can be found in Fig. 3. Intuitively, there are two main differences from the loss function (5). The first is the additional distortion term to optimize the transmission distortion $d(\mathbf{x},\mathbf{\hat{x}_{\text{H}}})$. The second is the scaling parameter $\alpha$ to control the relation between the entropy of the latent $\mathbf{z}_{l}$ and the transmission rate for $\mathbf{s}_{l}$.
To flexibly adjust the CBR for different images, a natural choice for reference is the prior of latent variable $\mathbf{z}_{l}$ that is correlated to the bit length after entropy coding, where we assume the entropy of $\mathbf{z}_{l}$ is positively related to the entropy of $\bm{\mu}_{l}$. Thus, the CBR for encoding $\bm{\mu}_{l}$ is designed to be proportional to the prior $p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$. Moreover, noting that the channel bandwidth is determined by reducing the number of the channel symbols, we achieve rate adjustment by masking the channel symbol vector according to the transmission rate $\alpha p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$, which will be introduced in the following sections.
Figure 5: The illustration of the process of the JSCC encoder and JSCC decoder for transmitting latent representation $\bm{\mu}_{l}$.
III-D HJSCC with Feedback
Building upon previous designs, we further investigate the JSCC scenario with feedback link from the receiver to the transmitter. Although feedback signal do not increase channel capacity, they simplify the coding mechanism, leading to better performance gains. In this scenario, image transmission is divided into multiple phases, allowing the transmitter to use the received channel symbol vectors from previous phases when encoding channel symbols in the current phase.
As shown in Fig. 4, we formulate our HJSCC with feedback by viewing the received symbols vector $\mathbf{\tilde{s}}_{l}$ as the latent variable. In this case, the loss function for training HJSCC with feedback can be written as
@@ III-D HJSCC with Feedback
We aim to achieve rate-adaptive transmission for HJSCC in the presence of feedback link. Similar to the scenario without feedback, we are able to take the prior information $p_{\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l})$ as a rate indicator. However, the calculation of this prior depends on the values of $\mathbf{\tilde{s}}_{l}$, which is unknown before transmission. As shown in Fig. 4, we address this by proposing a new prior $p_{\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l})$ as a substitution, where $\mathbf{z}_{l}$ is sampled from a uniform distribution centered on $\bm{\mu}_{l}$. This is inspired by the fact that $p_{\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l})$ is positively related to $p_{\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l})$, and thus can be employed as the rate term when training the model. As a result, the loss function for training the HJSCC with feedback can be written as

### v5

@@ Abstract
In this paper, we introduce an innovative hierarchical joint source-channel coding (HJSCC) framework for image transmission, utilizing a hierarchical variational autoencoder (VAE). Our approach leverages a combination of bottom-up and top-down paths at the transmitter to autoregressively generate multiple hierarchical representations of the original image. These representations are then directly mapped to channel symbols for transmission by the JSCC encoder. We extend this framework to scenarios with a feedback link, modeling transmission over a noisy channel as a probabilistic sampling process and deriving a novel generative formulation for JSCC with feedback. Compared with existing approaches, our proposed HJSCC provides enhanced adaptability by dynamically adjusting transmission bandwidth, encoding these representations into varying amounts of channel symbols. Additionally, we introduce a rate attention module to guide the JSCC encoder in optimizing its encoding strategy based on prior information. Extensive experiments on images of varying resolutions demonstrate that our proposed model outperforms existing baselines in rate-distortion performance and maintains robustness against channel noise.
Introduction
@@
To meet the transmission requirements of heavy data traffic in future sixth-generation (6G) networks, wireless edge devices need to be equipped with higher transmission efficiency. Most contemporary systems employ a two-step strategy for data transmission: first, the raw data is compressed using a source codec, such as JPEG (Wallace 1992) and BPG (Lainema et al. 2012). Then, the encoded bits are protected with redundancy introduced by a carefully designed channel codec, such as LDPC and Polar codes (Arikan 2009). However, in many practical applications, the bit length is generally finite, making it impossible to guarantee optimality. In this context, joint source-channel coding (JSCC) has emerged as a potential solution, offering higher coding gains than the traditional separation-based coding paradigm.
@@
With the revolutionary progress of deep learning in various fields, such as image compression (Ballé et al. 2018; He et al. 2021; Xu et al. 2022b) and generative models (Kingma and Welling 2022; Razavi, van den Oord, and Vinyals 2019), a novel design paradigm for JSCC, called learned image transmission (LIT), has been conceived by formulating the communication pipeline as an end-to-end deep learning model (Bourtsoulatze et al. 2019; Zhang et al. 2024; Sun et al. 2023). Specifically, these methods leverage powerful neural networks to implement the encoding and decoding processes. In this approach, the whole system is viewed as an autoencoder (AE), which can be jointly learned in a data-driven manner. A notable method proposed by (Bourtsoulatze et al. 2019) employed CNNs to construct the source and channel codecs for wireless image transmission, achieving great performance by mapping the input image directly into channel symbols. Moreover, Kurka and Gündüz; Wu et al. investigate the JSCC with feedback. In this context, the transmission of these representations is divided into multiple phases, with the transmitter receiving the channel symbol vector after each phase, which simplifies the encoding process and improves the overall performance.
@@
Beyond deterministic AEs, some studies have employed variational autoencoders (VAE) to design JSCC systems (Choi et al. 2019; Bo et al. 2024; Hu et al. 2023), where the channel symbols are generated through sampling. Part of these VAE-based methods show superior performance compared to deterministic AE-based methods, particularly under severe channel conditions. Though VAE-based methods have demonstrated remarkable performance, they experience significant performance degradation on high-resolution images. Furthermore, most existing methods only support fixed-rate coding, which contrasts with emerging works on transform coding-based image compression (Ballé et al. 2018), where the compression rate for each image is determined by the estimated entropy of its feature representation and varies with different samples. Consequently, these methods are less flexible and adaptive, potentially leading to performance penalties.
@@
In this work, we aim to overcome the limitations of previous methods while enhancing performance. Specifically, we develop a hierarchical JSCC (HJSCC) framework based on a powerful hierarchical VAE architecture (Child 2020). Our transmitter employs both bottom-up and top-down paths to autoregressively generate multiple hierarchical representations of the original image. These representations are then mapped to channel symbols using multiple JSCC encoder blocks. Building upon this, we further explore the application of HJSCC in a classical scenario where a feedback link exists. By modeling transmission over a noisy channel as a probabilistic sampling process, we derive a novel generative formulation for JSCC with feedback, which achieves significantly better performance than most existing advanced schemes. While there have been attempts at variable-rate transmission (Dai et al. 2022; Song et al. 2023; Zhang et al. 2023; Yang and Kim 2022) in the realm of JSCC without feedback, the problem of rate-adaptive design for JSCC with feedback remains underexplored. Unlike existing works (Kurka and Gündüz 2020; Wu et al. 2024a; Li et al. 2024; Jiang et al. 2022), we leverage the prior distribution (which characterizes the entropy information) of each representation to generate masks that control the number of symbols for each representation. This approach allows us to dynamically adjust the transmission rate. Additionally, we introduce a rate attention module to guide the JSCC encoder in adjusting the encoding strategy according to its prior information.
@@
VAE can be employed as deep generative models capable of generating high-dimensional data based on a low-dimensional latent space, and is a variant of autoencoder (Sø nderby et al. 2016; Cemgil et al. 2020). By sampling from the learned latent space distribution and passing these samples through the decoder network, VAE can generate new data points that resemble the training data. However, the original VAE is known to perform worse than many other generative models, particularly when applied to high-resolution images. (Vahdat and Kautz 2020; Child 2020; Yuhta Takida 2024) addressed this by proposing a deep hierarchical VAE, where the latent variable is divided into several disjoint groups, achieving significantly better performance than standard VAEs. More recently, VAE has been applied to compression tasks (Duan et al. 2023; Lu et al. 2024; Townsend, Bird, and Barber 2019; Kingma, Abbeel, and Ho 2019).
@@
Unlike the separation-based design described above, recent studies have delved into the utilization of AE and its variants, e.g., VAE to design wireless image transmission systems, resulting in a number of efficient methods (Bourtsoulatze et al. 2019; Xu et al. 2022a; Zhang et al. 2024; Sun et al. 2023; Xu et al. 2022a; Yang et al. 2024; Saidutta, Abdi, and Fekri 2019). In particular, (Bourtsoulatze et al. 2019; Wu et al. 2024b) and (Yang et al. 2024) conceived of using neural networks to simultaneously finish the source encoding/decoding and channel coding/decoding, with the goal of jointly optimizing the entire system to maximize PSNR. The VAE-based methods adopted probabilistic modeling, wherein the encoding process is characterized as a stochastic procedure. In these systems, channel symbols are generated by sampling from the probability distribution conditioned on the input image (Bo et al. 2024; Saidutta, Abdi, and Fekri 2019). These approaches have demonstrated superior performance, especially under severe transmission conditions.
Proposed Methods
Background
@@
VAE and Hierarchical VAE
As stated in (Cemgil et al. 2020), VAE is a stochastic variational inference scheme that can be applied to various intelligent tasks, such as recognition, denoising, and generation. As shown in Fig. 1(a), to formulate a vision-related model, we typically start with the following premises. Let $\mathbf{x}$ denote an image intensity vector, which is drawn from a dataset $\mathcal{X}$ with distribution $p_{\mathbf{x}}$. Another variable is the latent variable $\mathbf{z}$, with a prior $p_{\mathbf{z}}$. The main target of a VAE is to learn a generative model (decoder) $p_{\mathbf{x}|\mathbf{z}}$ for sampling, and a posterior density model (encoder) $q_{\mathbf{z}|\mathbf{x}}$ for variational inference. The objective for learning a VAE model can be formulated as minimizing the (variational) upper bound on the marginal likelihood of a batch of data points, as given by
$\mathcal{L}\left(\boldsymbol{\theta},\boldsymbol{\phi}\right)=\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\big{[}D_{KL}\left(q_{\mathbf{z}|\mathbf{x}}\|p_{\mathbf{z}}\right)-\log p_{\mathbf{x}|\mathbf{z}}\big{]},$ (1)
where $D_{KL}\left(q_{\mathbf{z}|\mathbf{x}}\|p_{\mathbf{z}}\right)=\log\big{(}q_{\mathbf{z}|\mathbf{x}}/p_{\mathbf{z}}\big{)}$, $\boldsymbol{\theta}$ and $\boldsymbol{\phi}$ represent the parameters of the encoder $q_{\mathbf{z}|\mathbf{x}}$ and decoder $p_{\mathbf{x}|\mathbf{z}}$, respectively.
@@
$\mathcal{L}\left(\boldsymbol{\theta},\boldsymbol{\phi}\right)=$ (2)
$\quad\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\left[\sum_{l=1}^{L}D_{KL}\left(q_{\mathbf{z}_{l}|\mathbf{z}_{<l},\mathbf{x}}\|p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}\right)\!-\!\log p_{\mathbf{x}|\mathbf{z}}\right],$
@@
Proposed HJSCC
System Overview
@@
Here, we aim to give a brief overview of deep learning-based JSCC. In particular, the model of a typical JSCC is shown in Fig. 2. The transmitter employs a JSCC encoder to map the input image $\mathbf{x}\in\mathbb{R}^{N}$ directly into channel symbol vector $\mathbf{s}\in\mathbb{C}^{K}$ for transmission, where $N$ denotes the number of pixels and $K$ represents the number of channel symbols. This process can be expressed as $\mathbf{s}=g_{e}(\mathbf{x};\boldsymbol{\theta}_{e})$, where $g_{e}$ signifies the encoding function and $\boldsymbol{\theta}_{e}$ represents its trainable parameters. Accounting for the limited transmission power, the transmitted signal should satisfy the power constraint $P$, implying that $\|\mathbf{s}\|^{2}_{2}/K\leq P$. Subsequently, the channel symbol vector is transmitted through the wireless channel, as given by $\mathbf{\tilde{s}}=\mathbf{s}+\mathbf{n}$, where $\mathbf{n}\sim\mathcal{N}(0,\sigma^{2}_{n}\mathbf{I})$ is additive white Gaussian noise (AWGN). At the receiver, the received noisy signal $\mathbf{\tilde{s}}$ is processed by a decoder $g_{d}$ to obtain the reconstructed image $\mathbf{\bar{x}}=g_{d}(\mathbf{\tilde{s}};\boldsymbol{\phi}_{d})$, with $\boldsymbol{\phi}_{d}$ denoting the trainable parameters of the JSCC decoder. The optimization objective of a deep learning-based JSCC system is to minimize the difference between $\mathbf{x}$ and $\mathbf{\bar{x}}$, and thus mean-square error (MSE) can be employed as the loss function. Furthermore, to evaluate the performance of a JSCC system and ensure fairness, we define the signal-to-noise ratio (SNR) as SNR =$\text{SNR }=10\log\frac{P}{\sigma^{2}}(\mathrm{dB})$, which characterizes the channel quality of the system. Then, we introduce the channel bandwidth ratio (CBR) to describe the transmission rate (overhead), which is expressed as $\text{CBR}=K/N$. Intuitively, CBR actually signifies the number of symbols for transmitting one pixel, and a higher CBR brings a higher overhead, while usually resulting in a better system performance.
@@
While previous studies have achieved excellent rate-distortion performance based on the framework depicted in Fig. 2, it is evident that the transmission rate is solely determined by the image resolution. This limitation can result in performance degradation in overall rate-distortion due to the inability to adaptively adjust the rate for each image (He et al. 2022). To address this issue, we propose our HJSCC framework, as illustrated in Fig. 3. Given an image $\mathbf{x}$, the bottom-up path generates a set of latent features, which are subsequently passed to the top-down path to autoregressively generate the latent representations $\boldsymbol{\mu}\triangleq\left\{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\ldots,\boldsymbol{\mu}_{L}\right\}$. These representations are then fed to a set of JSCC encoders $g_{e}=\left\{g^{1}_{e},g^{2}_{e},\ldots,g_{e}^{L}\right\}$, respectively. In this way, we are able to obtain a set of channel symbol vectors $\mathbf{s}\triangleq\left\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L}\right\}$, where $\mathbf{s}_{l}=g_{e}^{l}(\mathbf{z}_{l})$. Then, these channel symbol vectors are transmitted to the receiver through the wireless link, and the received symbol vectors are represented by $\mathbf{\tilde{s}}\triangleq\left\{\mathbf{\tilde{s}}_{1},\mathbf{\tilde{s}}_{2},\ldots,\mathbf{\tilde{s}}_{L}\right\}$. At the receiver, the noisy $\mathbf{\tilde{s}}$ undergoes processing by the JSCC decoder $g_{d}^{l}$, and we obtain $\boldsymbol{\tilde{\mu}}\triangleq\left\{\boldsymbol{\tilde{\mu}}_{1},\boldsymbol{\tilde{\mu}}_{2},\ldots,\boldsymbol{\tilde{\mu}}_{L}\right\}$. With $\boldsymbol{\tilde{\mu}}$ at hand, the receiver can reconstruct the image using the top-down path (decoder), and the reconstructed image is denoted by $\mathbf{\hat{x}}_{\text{H}}$. Our objective is to minimize the distortion between the transmitted image $\mathbf{x}$ and $\mathbf{\hat{x}}_{\text{H}}$.
@@
Training Objective Formulation
We aim to develop a rate-adaptive JSCC model capable of adjusting the transmission rate, CBR, based on the source content. To this end, we propose an inherited training strategy motivated by (Dai et al. 2022), where the objective of minimizing $d(\mathbf{x},\mathbf{\hat{x}}_{\text{H}})$ is guided by a learned image coder. Specifically, we have designs for the posteriors, priors, and training objectives as follows (Duan et al. 2023).
@@
where $\mathcal{U}$ denotes a uniform distribution centered on $\mu_{l}^{(i)}$, and $\mu_{l}^{(i)}$ is the $i$-th element of parameter $\boldsymbol{\mu}_{l}$, which is obtained from the $l$-th posterior branch.
Priors: The conditional prior distribution $p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$ is defined to be a Gaussian distribution convolved with a uniform distribution (Ballé et al. 2018):
@@
Training Loss: With the posteriors and priors defined above, the loss function (2) for an image compression model (Duan et al. 2023) can be expressed as
@@
$\mathcal{L}$ $=\mathbb{E}_{\mathbf{x}\sim p_{\mathbf{x}},\mathbf{z}\sim q_{\mathbf{z}|\mathbf{x}}}\big{[}$ (6)
$\quad\underbrace{\sum_{l=1}^{L}-\alpha\log p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}(\mathbf{z}_{l}|\mathbf{z}_{<l})}_{\text{transmission rate}}+\lambda\cdot(\underbrace{d(\mathbf{x},\mathbf{\hat{x}})+d(\mathbf{x},\mathbf{\hat{x}}_{\text{H}}))}_{\text{weighted distortion}}\big{]},$
where the calculation ways of each variable can be found in Fig. 3. Intuitively, there are two main differences from the loss function (5). The first is the additional distortion term to optimize the transmission distortion $d(\mathbf{x},\mathbf{\hat{x}_{\text{H}}})$, which helps to obtain a robust representation against noise. The second is the scaling parameter $\alpha$ to control the relation between the entropy of the latent $\mathbf{z}_{l}$ and the transmission rate for $\mathbf{s}_{l}$.
To flexibly adjust the CBR for different images, a natural choice for reference is the prior of latent variable $\mathbf{z}_{l}$ that is correlated to the bit length after entropy coding, where we assume the entropy of $\mathbf{z}_{l}$ is positively related to the entropy of $\boldsymbol{\mu}_{l}$. Thus, the CBR for encoding $\boldsymbol{\mu}_{l}$ is designed to be proportional to the prior $p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$. Moreover, noting that the channel bandwidth is determined by reducing the number of the channel symbols, we achieve rate adjustment by masking the channel symbol vector according to the transmission rate $\alpha p_{\mathbf{z}_{l}|\mathbf{z}_{<l}}$, which will be introduced in the following sections.
Figure 5: The illustration of the process of the JSCC encoder and JSCC decoder for transmitting latent representation $\boldsymbol{\mu}_{l}$.
HJSCC with Feedback
Building upon previous designs, we further investigate the JSCC scenario with feedback link from the receiver to the transmitter. In this scenario, image transmission is divided into multiple phases, allowing the transmitter to use the received channel symbol vectors from previous phases when encoding channel symbols in the current phase.
As shown in the left framework of Fig. 4, we formulate our HJSCC with feedback by viewing the received symbols vector $\mathbf{\tilde{s}}_{l}$ as the latent variable. In this case, the loss function for training HJSCC with feedback can be written as
@@
We aim to achieve rate-adaptive transmission for HJSCC in the presence of feedback link. Similar to the scenario without feedback, we are able to take the prior information $p_{\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l})$ as a rate indicator. However, the calculation of this prior depends on the values of $\mathbf{\tilde{s}}_{l}$, which is unknown before transmission. As shown in the right of Fig. 4, we address this by proposing a new prior $p_{\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l})$ as a substitution, where $\mathbf{z}_{l}$ is sampled from a uniform distribution centered on $\boldsymbol{\mu}_{l}$. This is inspired by the fact that $p_{\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{z}_{l}|\mathbf{\tilde{s}}_{<l})$ is positively related to $p_{\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l}}(\mathbf{\tilde{s}}_{l}|\mathbf{\tilde{s}}_{<l})$, and thus can be employed as the rate term when training the model. As a result, the loss function for training the HJSCC with feedback can be written as